This is week five of LSSTCam on-sky commissioning.
Clouds and colder temperatures have arrived at Cerro Pachón. During a period of in-dome testing, the team collected the most extensive set of in-dome calibration data to date. Measurements included wavelength scans for multiple filters using the Collimated Beam Projector (CBP). The team sampled the system optical throughput for multiple optical paths through the system by co-pointing the Simonyi telescope and CBP, illuminating specific sectors of the pupil. Wavelength scans were also acquired with no filter online, yielding measurements of the quantum efficiency for each individual detector. In addition, the team used the CBP to directly verify crosstalk matrices for LSSTCam by placing a spot on each amplifier and measuring crosstalk signals on adjacent amplifiers.
Calibration activities also included further commissioning of the flat-field system. For the broadband LED flats, two LED sources with emission in distinct wavelength ranges are used to fully cover each of the griz passbands. By using one LED at a time, the team made initial measurements of chromatic variations across the focal plane. Studies of the chromatic response will be expanded in the near future using monochromatic flats that systematically scan in wavelength.
As noted last week, the Camera focal plane cooling system experienced an anomaly, and the system was restarted. A preliminary environmental cause was identified by the team, along with mitigation options that are currently under study for implementation.
On-sky testing emphasized work to improve the consistency of the delivered image quality across a range of conditions. The team completed an initial verification of the updated open loop look-up tables and pointing model based on measurements from last week, and made progress optimizing the AOS closed loop configuration to reduce first-order spherical aberrations. By the end of the week, the team ran a series of tests of the AOS closed loop with all 50 degrees of freedom enabled for the first time. Other AOS tests involved running the AOS closed loop in “full array mode” by taking triplets of intra-focal, extra-focal, and in-focus images, similar to AOS commissioning with the LSST Commissioning Camera (LSSTComCam) in late 2024, but on a much larger scale. This computationally intensive operation solves the optical wavefront for the full array of 189 science sensors on the focal plane, and iteratively applies corrections to optimize delivered image quality across the field of view. Looking ahead, the team is now poised to begin a systematic study to characterize the thermal response of the M1M3 mirror figure and optimize control loops. In addition, the recent pointing model updates will facilitate testing with the guider sensors to evaluate higher temporal frequency contributions to the delivered image quality.
All images coming from LSSTCam have been processed at the US Data Facility (USDF). After dealing with some bottlenecks, Prompt Processing now processes each image in near real time. The USDF also supports Data Release Processing (DRP) campaigns for each night’s images, and selected observations from prior nights, with ever improving pipeline releases. USDF continues to optimize strategies for managing and moving data volumes within the embargo storage to accommodate the extended 30-day embargo period during on-sky engineering. USDF is well on the way to being a stable processing environment capable of handling the LSSTCam data deluge.
The COSMOS LSST Deep Drilling Field (DDF) was observed in the ugrizy bands with roughly 30 visits per band. These observations include the first on-sky images with the z and y filters. COSMOS is the first field to have full six-band photometry with LSSTCam. COSMOS is setting for the season. The four LSST DDFs located in the south Galactic cap are not currently visible, but will be accessible in a few months. The team also began a set of densely dithered observations in a moderately dense stellar field near to the Galactic plane to refine the camera distortion model and illumination correction. As part of Science Pipelines commissioning, the team will compare repeated measurements of the same stars appearing at many locations across the focal plane to evaluate the internal astrometric and photometric calibration.
On 20 May, the y filter was removed and the u filter was installed. The currently loaded filter set is ugriz as we enter dark time. The next new moon is around 26 May 2025.
See the recent announcement – Save the Date for Rubin First Look: 23 June 2025